A compound with the empirical formula of C2OH2. How much does the hydroxide portion weigh in regards to the entire compound?
1
Expert's answer
2020-12-18T08:41:11-0500

Empirical formula is "C_2OH_2"

The total molecular mass is

"(12.01*2)+(1*16.00)+(2*1.00)=42.02g\/mol"

The percentage hyroxide portion is

"12+2=14g\/mol"

"\dfrac{14g/mol}{42.02g/mol}×100=33\%"